OpenStudy (anonymous):

find the zeros for each function y=x3-5x^2+16x-80

Directrix (directrix):

0= x^3 - 5x^2 + 16x - 80 0 = (x - 5) (x ^ 2 + 16) 0- x - 5 OR (x + 4i) ( x - 4 i) = 0 x = 5 or x = -4i or x = 4i

OpenStudy (anonymous):

thank u thats what i had but i kep gettin the first part wrong

Directrix (directrix):

That's great.

Directrix (directrix):

The graph is cool.. Take a look here: http://www.wolframalpha.com/input/?i=x%5E3+-+5x%5E2+%2B+16x+-+80+%3D+y